40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed Mar 20 22:18:28 2013, tara, Notes, RefCav, estimated beat frequency 
    Reply  Fri Apr 5 13:56:01 2013, tara, Notes, RefCav, estimated beat frequency 
       Reply  Sat Apr 6 09:32:35 2013, rana, HowTo, TempCtrl, estimated beat frequency 
          Reply  Mon Apr 8 15:15:06 2013, rana, HowTo, TempCtrl, estimated beat frequency 
             Reply  Sun Apr 14 16:51:21 2013, rana, HowTo, TempCtrl, estimated beat frequency 
Message ID: 1142     Entry time: Fri Apr 5 13:56:01 2013     In reply to: 1124     Reply to this: 1143
Author: tara 
Type: Notes 
Category: RefCav 
Subject: estimated beat frequency 

Note:A test to check which cavity needs to be heated up.

1) when heaters (on the shields) are off, C3:PSL-VAC_CHMBRTMP =31.2

  • CAV1 is locked @ (slow out Coarse/fine) = 774/500
  • CAV2                 @ 769/500

2 when heater on CAV2(4V), C3:PSL-VAC_CHMBRTMP = 31.2

  • CAV1             @ 775/500
  • CAV2              @ 756/500

So to bring both cavity to be resonant at the same time, the heater on cav1 should be on.

I'm not sure if the thermometers on the shields are working or not, I'll check them.

Quote:

At current temperature, the estimated beat frequency will be ~ 60-100MHz. This is not so bad, since we can use 1811 to measure the beat signal and use PLL to extract the beat noise.

 

We will need to use thermal expansion to tune the beat frequency. So, as a start, I try to figure out the beat frequency, and how much we have to heat up the cavity. The heaters on each cavity is off, only the heater around the vac chamber is on (but the servo is off). 

Right now we have one laser locked to one cavity, but the beam path to the first cavity has a beam splitter that we can borrow the beam and direct it to the 2nd cavity. I realigned the beam to have both beams into both cavities. By adjusting the temperature control on the NPRO (slow signal), I can bring the beam to resonant in each cavity.

1st cav is resonant @ (334/398) and (150/398). The numbers correspond to coarse and fine knobs of the slow feedback to the laser.

2nd cav is resonant @ (154/398). (I'll come up with a better name to call the cavities)

The FSR is 4.07 GHz (for 1.45" long cavity). This means 334-154 = 184 clicks on the coarse knob equals to 4.07 GHz, or 22MHz per coarse click. Both cavities resonant at ~3-5 clicks apart. So the beat frequency is ~ 60-100MHz. This is quite good, at least we are not close to half FSR apart. The power required to tune the cavity length should not be that high.

 

The next thing to do is try to see which cavity we need to heat up in order to bring both cavities resonant frequency closer together.

 

ELOG V3.1.3-